# Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles

### v1

@@ Abstract
In Fig. 3 we plot the connected third‐order correlation $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ using the approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles) for an atomic ensemble with $\beta=1\%$ and $5\%$. The value $\beta=1\%$ represents the approximate value from past experiments [3, 6, 21], while $\beta=5\%$ is what could be obtained if the atoms are trapped more closely to the fibre and/or if the light-matter coupling is enhanced, for example, with slow light using fibre gratings. The result of $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ is shown in Jacobi coordinates $R=\frac{x_{1}+x_{2}+x_{3}}{\sqrt{3}},\quad\eta=\frac{x_{1}-x_{2}}{\sqrt{2}},\quad\zeta=\sqrt{\tfrac{2}{3}}\left(\frac{x_{1}+x_{2}}{2}-x_{3}\right)$ with $R=0$ fixed by translational invariance. In the $\eta$$\zeta$ plane, $g_{c}^{(3)}$ exhibits an apparent six-fold symmetry: the underlying three-fold bosonic permutation symmetry is doubled because $g_{c}^{(3)}$ does not distinguish whether a photon pair arrives before or after the lone photon. As the optical depth increases, $g_{c}^{(3)}$ evolves from being everywhere negative, through a regime of uniform positivity, to a characteristic pattern with a bright central peak surrounded by six negative “outer legs.” Under approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles), $g_{c}^{(3)}(x_{1},x_{2},x_{3})=2\phi_{3}(x_{1},x_{2},x_{3})/t_{0}^{3M}+2[\phi_{2}(x_{2},x_{1})\phi_{2}(x_{1},x_{3})+\phi_{2}(x_{1},x_{2})\phi_{2}(x_{2},x_{3})+\phi_{2}(x_{1},x_{3})\phi_{2}(x_{3},x_{2})]/t_{0}^{4M}+\mathcal{O}(\beta^{3})$. The low-OD behaviour is dominated by the amplitude of the 3-vertex diagram transport in $\phi_{3}$ because it has the lowest power in $\beta$. This amplitude carries a negative sign in position space. As OD increases, the amplitude of the 4-vertex diagram contribution to $\phi_{3}$ (with a positive sign) becomes significant. Meanwhile, the second term—being the product of two negative $\phi_{2}$ amplitudes—adds a net positive contribution. The competition between various transport mechanism at different OD explain the observed sign reversal and the emergence of the six-lobed pattern at high optical depth.
When the number of the atoms is small $M\leq 11$, $g_{c}^{(3)}$ can be numerically calculated using the quantum regression theorem (QRT) and the cascaded master equation. The comparison between $g_{c}^{(3)}$ obtained by our theory and simulation is given in Fig. 6. To benchmark the accuracy of our perturbative calculation, we computed $g_{c}^{(3)}(x_{1},x_{2},0)$ on a $50\times 50$ grid ($x_{1},x_{2}\in[0,5/\Gamma_{\rm tot}]$) using both our diagrammatic analytic calculation and QRT, yielding matrices $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. We calculate the relative error between theory and QRT using $\epsilon=\|\mathcal{M}-\mathcal{M}^{\prime}\|_{F}/\|\mathcal{M}\|_{F}$, where $\|\cdot\|_{F}$ is the Frobenius norm. For an array with $\beta=5\%$, $\epsilon<2.0\%$ and $\epsilon<6.2\%$ for drive powers $P_{\rm in}=0.02\Gamma_{\rm tot}$ and $P_{\rm in}=0.06\Gamma_{\rm tot}$ respectively. These results confirm the quantitative accuracy of our loop-level perturbative method at low OD ($\text{OD}\leq 0.4$). We also see from Fig. 6 that even though the value of $g_{c}^{(3)}$ diminishes at higher power $P_{\rm in}\geq 0.4\Gamma_{\rm tot}$, the simulations still qualitatively agree with the theory at $\mathrm{OD}>1$. This is because: for the approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles) under the weak-drive condition, we only considered three-photon diagrams for $g^{(3)}$ and two-photon diagrams for $g^{(2)}$. When the drive power is further increased to $P_{\rm in}/\Gamma_{\rm tot}=\mathcal{O}(1)$, $P_{\rm in}/\Gamma_{\rm tot}$ is no longer a good expansion parameter, and thus the photon transport diagrams with more photons, such as Figs. 4-5, should be taken into consideration.
To observe correlated/anti-correlated photon triples, $P_{\rm in}$ and OD should be chosen such that $|g_{c}^{(3)}|$ is sufficiently large and the count rate of triples should be in the detectable range. We estimate the experimental counting rate $\mathcal{S}$ of arriving within the window time $3/\Gamma_{\rm tot}$ as
$\mathcal{S}=\int_{0}^{3/\Gamma_{\text{tot}}}dt_{1}dt_{2}\,|G_{c}^{(3)}(t_{1},t_{2},0)|.$ (7)
Here, $G_{c}^{(3)}(t_{1},t_{2},t_{3})=g_{c}^{(3)}\prod_{j=1}^{3}\expectationvalue{\hat{a}^{\dagger}(t_{j})\hat{a}(t_{j})}$. We consider parameters $\Gamma_{\rm tot}=2\pi\times 5$ MHz and fixed driving strength $P_{\rm in}=0.06\Gamma_{\rm tot}$. Forg an OD range with detectable three-photon correlations such that $|g_{c}^{(3)}|\geq 0.1$, our estimation shows that the correlated-triple-photon count rate is 5 Hz and 2 Hz for $\beta=5\%,M=8$ and $\beta=1\%,M=85$, respectively. For larger powers, $\mathcal{S}\propto P_{\rm in}^{3}$, and thus $\cal S$ increases rapidly. This comes only at the expense of a slightly reduced $|g_{c}^{(3)}|$ as shown in the simulations Fig. 6. This is because most correction diagrams, shown in Figs. 4-5, decay exponentially with OD. Therefore, experiments can adjust driving strengths to balance count rates and $|g_{c}^{(3)}|$.
@@
Conclusion.—We have developed a diagrammatic scattering theory to calculate three-photon transport through weakly coupled atomic ensembles. Our formalism provides analytic expression of the leading-order processes that contribute to non-separable three-photon correlations. We have used our new approach to predict the presence of non-Gaussian correlations in the transmitted field of coherently driven atomic ensembles. We anticipate our results will pave the way for experimental demonstration of sources of non-Gaussian photon states using atoms coupled to nanofibres.
@@
Appendix A End Matter
A.1 The Diagrams for the non-Perturbative Drive Regime
In the weak-driving regime, where $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta)$, the order estimation combined with the approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles) indicates that it suffices to consider single-, two-, and three-photon diagrams for evaluating the correlators $\expectationvalue{:\prod_{j=1}^{n}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):}$ with $n=1,2,3$, respectively, in order to obtain $g_{c}^{(3)}(x_{1},x_{2},x_{3})$. When the drive power increases to $P_{\rm in}/\Gamma_{\rm tot}=\mathcal{O}(1)$, the diagrams with additional photons must be included in each $\expectationvalue{:\prod_{j=1}^{n}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):}$ to obtain an accurate perturbative result. This, however, becomes formidable due to the rapid growth in combinatorial complexity of diagrams.
Numerical simulations in Fig. 6 reveal that $g_{c}^{(3)}$ exhibits markedly different temporal patterns under weak and strong driving at small OD, but the two regimes appear similar at large OD. This suggests that the contribution from the temporally dependent part of additional-photon diagrams may decay rapidly with OD. To illustrate this, we list below several representative terms from the diagrams with more photons and show their scaling with various powers of $t_{0}^{M}$. In the figures on the left, the coordinate $x_{j}$ attached to the end of a photon-transport line denotes the outgoing photons contracted with the $\hat{a}(x_{j})$ or $\hat{a}^{\dagger}(x_{j})$ operator in the definition of the correlators. The cross attached at the end of the photon line indicates that the photon is lost at some atom in the array. These diagrams form part of the total sum represented by the figures on the right.
First, we examine the additional diagrams contributing to $\langle\hat{a}^{\dagger}(x)\hat{a}(x)\rangle$. Owing to the translational invariance of the steady state, these diagrams—like the output power itself—are independent of $x$. The representative term shown on the left in Fig. 4a scales as $t_{0}^{4M}$. In many cases, the scaling with the exponential decay factor $t_{0}^{M}$ can be inferred without performing the full, laborious diagrammatic evaluation of the expectation value. Instead, one can exploit the connectedness of the diagrams and momentum conservation in photon-photon interactions. For example, in Fig. 4a, the bra diagram contributes a factor $t_{0}^{2M}$. When evaluating the expectation value, the bottom photons in the bra and ket contract, enforcing the same outgoing zero momentum. Momentum conservation in the photon-photon interaction within the ket then requires the top photon in the ket to also have zero momentum. Since both photons have zero momentum before and after the interaction, the amplitude is $\sim t_{0}^{2M}$. From an explicit calculation for Fig. 4b (not shown here) we have seen that it does not exhibit explicit $t_{0}^{M}$ scaling, implying that its contribution to $g_{c}^{(3)}$ persists even as OD increases. However, because it enters $\langle\hat{a}^{\dagger}(x)\hat{a}(x)\rangle$ in the denominator of $g^{(3)}$ and $g^{(2)}$ as a constant, it rescales the overall magnitude of the correlation functions with increasing $P_{\rm in}$, without altering their temporal dependence.
$x$ $x$ $\subset$ $\hat{a}^{\dagger}(x)\hat{a}(x)$
(a) $\mathcal{O}(\beta)$
$x$ $x$ $\subset$ $\hat{a}^{\dagger}(x)\hat{a}(x)$
(b) $\mathcal{O}(\beta^{2})$
Figure 4: The representative additional term to the output power $\expectationvalue{\hat{a}^{\dagger}(x)\hat{a}(x)}$.
The terms which depends on the coordinate $x_{1}$ and $x_{2}$ appears in the additional diagram for $\expectationvalue{\hat{a}^{\dagger}(x_{1})\hat{a}^{\dagger}(x_{2})\hat{a}(x_{2})\hat{a}(x_{1})}$ and for $\expectationvalue{\hat{a}^{\dagger}(x_{1})\hat{a}^{\dagger}(x_{2})\hat{a}^{\dagger}(x_{3})\hat{a}(x_{3})\hat{a}(x_{2})\hat{a}(x_{1})}$. However, by following a similar analysis to that in the previous paragraph, we see that both the representative terms in Fig. 5a and 5b scale with $t_{0}^{4M}$. Meanwhile, both the representative terms in Fig. 5c and 5d scale with $t_{0}^{6M}$. One could enumerate all the other temporal dependent terms from other additional diagrams at $\mathcal{O}(\beta^{2})$ to see that they all decay at least as fast as $t_{0}^{4M}$. This rapid decay of the temporal dependent terms explains why $g_{c}^{(3)}$ exhibits similar temporal patterns at large OD when the system is under weak and strong drives.
$x_{1}$$x_{2}$ $x_{2}$$x_{1}$ $\subset$ $\prod_{j=1}^{2}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):$
(a) $\mathcal{O}(\beta^{2})$
$x_{1}$$x_{2}$ $x_{2}$$x_{1}$ $\subset$ $\prod_{j=1}^{2}\hat{a}^{\dagger}(x_{j})\hat{a}^{\dagger}(x_{j}):$
(b) $\mathcal{O}(\beta^{2})$
$x_{1}$$x_{2}$$x_{3}$ $x_{2}$$x_{1}$$x_{3}$ $\subset$ $\prod_{j=1}^{3}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):$
(c) $\mathcal{O}(\beta^{2})$
$x_{1}$$x_{2}$$x_{3}$ $x_{2}$$x_{1}$$x_{3}$ $\subset$ $\prod_{j=1}^{3}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):$
(d) $\mathcal{O}(\beta^{2})$
Figure 5: The representative additional term to $\expectationvalue{\hat{a}^{\dagger}(x_{1})\hat{a}^{\dagger}(x_{2})\hat{a}(x_{2})\hat{a}(x_{1})}$ in (a) and (b). These representative terms are proportional to $t_{0}^{4M}$, in the diagrams with additional photons. The representative additional term to $\expectationvalue{\hat{a}^{\dagger}(x_{1})\hat{a}^{\dagger}(x_{2})\hat{a}^{\dagger}(x_{3})\hat{a}(x_{3})\hat{a}(x_{2})\hat{a}(x_{1})}$ in (c) and (d). These terms are nothing but the additional term for $\expectationvalue{\hat{a}^{\dagger}(x_{1})\hat{a}^{\dagger}(x_{2})\hat{a}(x_{2})\hat{a}(x_{1})}$ multiplied by $P_{\rm in}t_{0}^{2M}$ because of the appearance of one more pair of contracted independent scattering photons at the bottom.
Theory $P_{\rm in}=0.02\Gamma_{\rm tot}$ $P_{\rm in}=0.06\Gamma_{\rm tot}$ $P_{\rm in}=0.1\Gamma_{\rm tot}$ $P_{\rm in}=0.2\Gamma_{\rm tot}$ $P_{\rm in}=0.4\Gamma_{\rm tot}$ $P_{\rm in}=0.8\Gamma_{\rm tot}$
$M=2$ $\text{OD}=0.4$
$M=4$ $\text{OD}=0.8$
$M=6$ $\text{OD}=1.2$
$M=8$ $\text{OD}=1.6$
$M=11$ $\text{OD}=2.2$
($a$)
Figure 6: $g_{c}^{(3)}$ for the atomic ensemble-fibre system with coupling strength $\beta=5\%$ computed by our diagrammatic approach (the first column) and by simulation by using cascaded master equation with various driving power (the second to the last column). The orange and red contour line shows the region where $|g_{c}^{(3)}|$ greater than 0.05 and 0.1, respectively.

### v2

@@ Abstract
In Fig. 3 we plot the connected third‐order correlation $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ using the approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles) for an atomic ensemble with $\beta=1\%$ and $5\%$. The value $\beta=1\%$ represents the approximate value from past experiments [Prasad2020, Mahmoodian2021, Mahmoodian2023], while $\beta=5\%$ is what could be obtained if the atoms are trapped more closely to the fibre and/or if the light-matter coupling is enhanced, for example, with slow light using fibre gratings. The result of $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ is shown in Jacobi coordinates $R=\frac{x_{1}+x_{2}+x_{3}}{\sqrt{3}},\quad\eta=\frac{x_{1}-x_{2}}{\sqrt{2}},\quad\zeta=\sqrt{\tfrac{2}{3}}\left(\frac{x_{1}+x_{2}}{2}-x_{3}\right)$ with $R=0$ fixed by translational invariance. In the $\eta$$\zeta$ plane, the value of $g_{c}^{(3)}$ at the origin $(\eta,\zeta)=(0,0)$ corresponds to a three-photon coincidence $x_{1}=x_{2}=x_{3}$, i.e., enhanced (positive) or suppressed (negative) simultaneous detection of all three photons. Three axes $\eta=0$, $\zeta=\sqrt{3}/3\eta$ and $\zeta=-\sqrt{3}/3\eta$ corresponding to two-photon coincidence events with the third photon separated (except at their intersection (0,0)). $g_{c}^{(3)}$ is symmetric with respect of the three axes because of bosonic permutation symmetry. Furthermore, this three-fold symmetry is doubled to be a six-fold symmetry because $g_{c}^{(3)}$ does not distinguish whether a photon pair arrives before or after the lone photon.
As the optical depth increases, $g_{c}^{(3)}$ evolves from being everywhere negative, through a regime of uniform positivity, to a characteristic pattern with a bright central peak surrounded by six negative “outer legs.” Under approximation (Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles), $g_{c}^{(3)}(x_{1},x_{2},x_{3})=2\phi_{3}(x_{1},x_{2},x_{3})/t_{0}^{3M}+2[\sum_{(i,j,k)\in A}\phi_{2}(x_{i},x_{j})\phi_{2}(x_{j},x_{k})]/t_{0}^{4M}+\mathcal{O}(\beta^{3})$, where $A=\{(1,2,3),(2,1,3),(1,3,2)\}$. In contrast, if we hypothesize that the outgoing state has only Gaussian field correlation, Isserlis’s theorem gives $\widetilde{g}_{c}^{(3)}(x_{1},x_{2},x_{3})=2[\sum_{(i,j,k)\in A}\phi_{2}(x_{i},x_{j})\phi_{2}(x_{j},x_{k})]/t_{0}^{4M}+\mathcal{O}(\beta^{3})$ [Cardin2024, WDM2025]. Hence, $g_{c}^{(3)}$ becomes the sum of two parts: (i) the amplitude of correlated three-photon wavefunction $\phi_{3}$ (included in the first term and the high-order terms collected in $\mathcal{O}(\beta^{3})$), (ii) the product amplitude of correlated two-photon wavefunctions $\phi_{2}\phi_{2}$. Non-Gaussian intensity correlation therefore have contributions from terms with (i) non-Gaussian and (ii) Gaussian field correlations.
In the following, we explain the spatial pattern in Fig. 3 in terms of the relative amplitude of various transport processes in Fig. 2 and two-photon interaction. From the formulae of Weyl transformation, we see that without photon loss, a $n$-photon interaction at an emitter multiplies a factor of $\beta^{n}$ (nn=2,3) on the corresponding wavefunction amplitude. At low OD $4\beta M=0.6$, $\phi_{3}$ is dominated by the scattering amplitude of the process, whose sign is also negative, depicted in Fig. 2a where three-photon interaction only happens once during transport, having a factor of $\beta^{3}$ and $\phi_{2}$ is dominated by a single two-photon interaction transport having a factor of $\beta^{2}$. The $\phi_{2}\phi_{2}$ part of $g_{c}^{(3)}$ is negligible at low OD plotted in Fig. 3(a),(b). As the optical depth increases, e.g., to $4\beta M=2$, more atom sites are available to accommodate photon-photon interactions. Sequential photon-photon interactions in the transports results in the accumulation of the amplitude of processes with positive sign in both the $\phi_{3}$ and $\phi_{2}\phi_{2}$. In $\phi_{3}$, the process involving two successive two-photon interactions (Fig. 2b) grows in importance because it is enhanced by the combinatorial factor $\binom{M}{2}$ counting the choice of two emitter sites. At the same time, the magnitude of $\phi_{2}$ itself becomes larger since the leading single two-photon interaction contribution is enhanced by a combinatorial factor $M$ from the $M$ possible sites at which the two-photon interaction can occur. At large OD $4\beta M=4$, $\phi_{2}\phi_{2}$ dominates as the $\phi_{3}$ term is weighted by one more factor of $t_{0}^{M}$ exponentially decaying with OD. Here, $g_{c}^{(3)}$ becomes approximately proportional to $\phi_{2}\phi_{2}$. Its spatial pattern in Fig. 3 (e),(f) can be explain solely from the position space behavior of $\phi_{2}$: $\phi_{2}$ is negative when two photons are nearly coincident and positive when they are at moderate separation. Hence, the product $\phi_{2}(x_{i},x_{j})\phi_{2}(x_{j},x_{k})$ is positive when three photon coincide and is negative when one is far from the other two. To visualize the violation from Isserlis’s theorem, we calculate $g_{c}^{(3)}-\widetilde{g}_{c}^{(3)}$ (in Fig. S3).
For a small number of atoms ($M\leq 11$), both $g_{c}^{(3)}$ and $\widetilde{g}_{c}^{(3)}$ can be computed numerically using the quantum regression theorem (QRT) together with the cascaded master equation. As shown in Fig.S4, although $g_{c}^{(3)}$ and $\widetilde{g}_{c}^{(3)}$ decrease at higher input power $P{\rm in}\geq 0.4\Gamma_{\rm tot}$, the simulations remain in qualitative agreement with the theory for $\mathrm{OD}>1$. The signal strength of $g_{c}^{(3)}$ is discussed in Supplementary Material and our accompanying work [WDM2025].
@@
Conclusion.—We have developed a diagrammatic scattering theory to calculate three-photon transport through weakly coupled atomic ensembles. Our formalism provides analytic expression of the leading-order processes that contribute to non-separable three-photon correlations. We have used our new approach to predict the presence of non-Gaussian intensity and field correlations in the transmitted field of coherently driven atomic ensembles. We anticipate our results will pave the way for experimental demonstration of sources of non-Gaussian photon states using atoms coupled to nanofibres.